Find the mobius transformation which maps the 

circle |𝑧| < 1 conformally onto |𝑤 − 1| < 1 and 

make the points z=0,1 correspond to w=1/2,0 

respectively. Is the transformation uniquely 

determined


1
Expert's answer
2021-03-11T04:21:21-0500

The Mobius transformation is generally given by

"f(z) = \frac{az+b}{cz+d}"

As "f(0)\neq \infty", we conclude that "d \neq 0" and thus we can choose "d=2" (as in Mobius transformation multiplying "a,b,c,d" by a same non-zero constant gives the same transformation). Knowing the images if "z=0, z=1" we find:

  • "\frac{b}{d}=\frac{1}{2}, b=1"
  • "a+b=0, a=-1"

Now suppose "z\in \mathbb{C}, |z|=1", then this tranformation should carry it onto "|w-1|=1", thus

"|f(z)-1|=|\frac{-z+1-(cz+2)}{cz+2}|=|\frac{-(c+1)z-1}{cz+2}|"

"|f(z)-1|=\frac{(1+(c+1)\cos \theta)^2+(c+1)^2\sin^2\theta }{(c\cos\theta +2)^2+c^2 \sin^2\theta}=1"

"1+(c+1)^2\cos^2\theta+2(c+1)\cos\theta +(c+1)^2\sin^2\theta =c^2\cos^2\theta+4c\cos\theta +4+c^2\sin^2\theta"

"1+(1+c)^2+2(c+1)\cos \theta = c^2 +4c\cos\theta+4"

"(1+2c)+2(1-c)\cos \theta =3"

"2(1-c)\cos \theta =2(1-c)"

As this should be satisfied for all "\theta \in [0;2\pi]", we have "c=1" and thus the Mobius transformation is

"f(z)=\frac{-z+1}{z+2}" and the couple "[-1:1:1:2]" is unique up to a multiplication by a scalar.